### v2

# Paired kernels and truncated Toeplitz operators

Abstract.
This paper considers paired operators in the context of the Lebesgue Hilbert space $L^{2}$ on the unit circle and its subspace, the Hardy space $H^{2}$. The kernels of such operators, together with their analytic projections, which are generalizations of Toeplitz kernels, are studied. Inclusion relations between such kernels are considered in detail, and the results are applied to describing the kernels of finite-rank asymmetric truncated Toeplitz operators.
Key words and phrases:
Paired operator, paired kernel, Toeplitz operator, nearly-invariant subspace, truncated Toeplitz operator, kernel
1991 Mathematics Subject Classification:
Primary 30H10; Secondary 47B35, 47B38
To Ilya Spitkovsky on his 70th birthday
1. Introduction
If $X$ is a Banach space, $P\in\mathcal{L}(X)$ is a projection, $Q=I-P$ is its complementary projection, and $A$ and $B$ are bounded operators on $X$, then we use the term paired operators for operators of the form
$S_{A,B}=AP+BQ.$
Paired operators appeared in the context of singular integral operators on $L^{2}(\mathbb{R})$ of the form
[…]
where $a$ and $b$ are bounded functions and $g$ is given. They were introduced by Shinbrot [25], where (1.1) was placed in an abstract Hilbert space setting and a general method of obtaining solutions to that equation was derived in that abstract form.
Here we shall take $X=L^{2}(\mathbb{T})$, where $\mathbb{T}$ is the unit circle, $P$ and $Q$ are the orthogonal projections from $L^{2}(\mathbb{T})$ onto the Hardy space $H^{2}_{+}:=H^{2}(\mathbb{D})$ and its orthogonal complement $H^{2}_{-}:=\overline{H^{2}_{0}}$, respectively (we write $P=P^{+}$ and $Q=P^{-}$), and $A,B$ are multiplication operators $A=M_{a}$ and $B=M_{b}$ with $a,b\in L^{\infty}:=L^{\infty}(\mathbb{T})$. Thus we write
(1.2) $S_{a,b}=aP^{+}+bP^{-}\quad\hbox{and}\quad\Sigma_{a,b}=P^{+}aI+P^{-}bI.$
In each case the pair $(a,b)$ is called the symbol pair. These operators have been studied by several mathematicians [5, 11, 13, 16, 17, 18, 22, 24, 26], often in the context of singular integral operators. If the operators $A$ and $B$ are invertible in $\mathcal{L}(X)$, then the paired operators are said to be of normal type [22], and this is the most studied case by far.
Recently, several new results were obtained for operators of the form (1.2) with $a,b\in L^{\infty}$ satisfying weaker conditions [5, 11]:
(1.3) $a\neq 0\hbox{ a.e.\ on }\mathbb{T},\quad b\neq 0\hbox{ a.e.\ on }\mathbb{T}$
and
(1.4) $\hbox{either}\quad a=b\quad\hbox{or}\quad a-b\neq 0\hbox{ a.e.\ on }\mathbb{T},$
which will be standing assumptions in this paper. If $a\neq b$ we say that $(a,b)$ is nondegenerate.
Here we focus on paired operators of the form $S_{a,b}$ and their kernels. Paired operators are closely connected with Toeplitz operators. Indeed, on the one hand they are dilations of Toeplitz operators; on the other hand, if $b\in\mathcal{G}L^{\infty}$ (that is, if $b$ is invertible in $L^{\infty}$), then $S_{a,b}$ is equivalent after extension to the Toeplitz operator $T_{a/b}$ [3], implying that those two types of operator share many properties such as invertibility and Fredholmness, and their kernels are isomorphic. (See, for instance, [1] for the definition and properties of operators that are equivalent after extension.)
Paired kernels, i.e., kernels of paired operators, are thus particularly connected with Toeplitz kernels, i.e., kernels of bounded Toeplitz operators. If $b\in\mathcal{G}L^{\infty}$, not only do we have
(1.5) $\mathop{\rm ker}\nolimits S_{a,b}\cong\mathop{\rm ker}\nolimits T_{a/b}$
but in fact
(1.6) $P^{+}\mathop{\rm ker}\nolimits S_{a,b}=\mathop{\rm ker}\nolimits T_{a/b}.$
The spaces of the form $P^{+}\mathop{\rm ker}\nolimits S_{a,b}$, denoted by ${\rm ker}^{+}S_{a,b}$, are called projected paired kernels. They appear in different forms: in the study of scalar-type kernels of block Toeplitz operators [10], in the study of nearly invariant subspaces for shift semigroups [21], in the characterization of the ranges of finite rank truncated Toeplitz operators (TTO) [4], and so on. In Section 2 we review some of the basic properties of such projected paired kernels, and describe their properties related to near invariance, a property that Toeplitz kernels are known to possess. From this point of view it is natural to ask whether further known properties of Toeplitz kernels can be extended, or have analogues, for paired kernels. In Section 3 we focus on the relations between Toeplitz kernels corresponding to Toeplitz operators whose symbols are related by a multiplication operator, such as the known relation
(1.7) $\mathop{\rm ker}\nolimits T_{\theta g}\subseteq\mathop{\rm ker}\nolimits T_{g}\quad\hbox{if }\theta\hbox{ is an inner function},$
where the inclusion is strict if $\theta$ is not a constant [6, 7]; moreover,
(1.8) $h_{+}\mathop{\rm ker}\nolimits T_{h_{+}g}\subseteq\mathop{\rm ker}\nolimits T_{g}\quad\hbox{if }h_{+}\in H^{\infty},$
where the inclusion is an equality if $h_{+}\in\mathcal{G}H^{\infty}$; also
(1.9) $\mathop{\rm ker}\nolimits T_{h_{-}g}\supseteq\mathop{\rm ker}\nolimits T_{g}\quad\hbox{if }h_{-}\in\overline{H^{\infty}},$
where the inclusion is an equality if $\overline{h_{-}}$ is outer.
The relations above can be roughly seen as meaning that, when we multiply the symbol of a Toeplitz operator by a given function, the kernel may become “smaller” or “larger” – which cannot be expressed otherwise unless we are dealing with finite-dimensional kernels.
We wish to study and express the relations between paired kernels, corresponding to symbol pairs related by multiplication operators, similarly. It is clear that
[…]
but inclusion relations analogous to (1.7)–(1.9), which in the finite-dimensional case express the fact that the dimension of one kernel is smaller than that of the other, are impossible to obtain because the only paired kernels contained in $\mathop{\rm ker}\nolimits S_{a,b}$ are $\{0\}$ and $\mathop{\rm ker}\nolimits S_{a,b}$ itself [5].
It can be shown, however, that under our standing assumptions (1.3) and (1.4) the projection $P^{+}$ maps $\mathop{\rm ker}\nolimits S_{a,b}$ bijectively onto $P^{+}\mathop{\rm ker}\nolimits S_{a,b}={\rm ker}^{+}S_{a,b}$, indeed $P^{+}\phi=\dfrac{b}{b-a}\phi$, given that (1.4) holds (see [11, Prop. 2.1] for further details).
Thus we study here the relations between the projected paired kernels with symbol pairs related by multiplication operators, for which one can obtain inclusions and equalities analogous to (1.7)–(1.9). These relations have also been sudied in [11], but we present them here in Section 3 for completeness, on the one hand, and also because several results are generalized. In Section 4 we apply them to study and describe the kernels of a class of asymmetric truncated Toeplitz operators (ATTO) of the form
(1.11) $A^{\theta,\alpha}_{\phi}:K_{\theta}\to K_{\alpha},\qquad A^{\theta,\alpha}_{\phi}=P_{\alpha}\phi P_{\theta}{}_{|K_{\theta}},$
where $\theta,\alpha$ are inner functions, $K_{\theta},K_{\alpha}$ the associated model spaces, and $P_{\theta},P_{\alpha}$ the orthogonal projections from $L^{2}$ onto these model spaces [8]. In addition we take $\phi\in L^{\infty}$ with
(1.12) $\phi=\overline{\theta}\frac{A_{1-}}{B_{1-}}-\alpha\frac{A_{2+}}{B_{2_{+}}},\qquad A_{1-},B_{1-}\in\overline{H^{\infty}},\quad A_{2+},B_{2+}\in H^{\infty},$
such that $(A_{2+},B_{2+})\in{\rm CP}^{+}$ and $(A_{1-},B_{1-})\in{\rm CP}^{-}$ (corona pairs, defined in Section 4), which are motivated by the study of ATTO of finite rank. This is done by resorting to a characterization of certain kernels of block Toeplitz operators that are called scalar-type kernels [10, Thm. 3.1] and using the results of the previous section: this involves establishing certain relations between the corona theorem and Hankel operators. It is shown, in particular, that $\mathop{\rm ker}\nolimits A^{\theta,\alpha}_{\phi}$ does not depend on $\alpha$ nor on the numerators $A_{1-}$ and $A_{2+}$ if $\dim K_{\alpha}$ is siufficiently large, which may be seen as rather a surprising result.
2. Projected paired kernels and near invariance
In this section we summarize some of the basic properties of projected paired kernels. A more detailed treatment is given in [11]. We remind the reader of our standing assumptions (1.3) and (1.4).
The projected paired kernel ${\rm ker}^{+}S_{a,b}=P^{+}\mathop{\rm ker}\nolimits S_{a,b}\subset H^{2}_{+}$ was introduced in Section 1, and analogously we define ${\rm ker}^{-}S_{a,b}=P^{-}\mathop{\rm ker}\nolimits S_{a,b}\subset H^{2}_{-}$. Note that $\phi_{-}\in{\rm ker}^{-}S_{a,b}$ if and only if for some $\phi_{+}\in H^{2}_{+}$ we have $a\phi_{+}+b\phi_{-}=0$; that is, $\overline{a}\overline{z}\overline{\phi_{+}}+\overline{b}\overline{z}\overline{\phi_{-}}=0$, and so we see that
(2.1) ${\rm ker}^{-}S_{a,b}=\overline{z}\overline{{\rm ker}^{+}S_{\overline{b},\overline{a}}}.$
We may rewrite this by saying that the antilinear mapping
(2.2) $T:{\rm ker}^{+}S_{\overline{b},\overline{a}}\to{\rm ker}^{-}S_{a,b},\qquad\psi_{+}\mapsto\overline{z}\overline{\psi_{+}}$
is a bijection.
There is likewise a bijection
(2.3) $M_{-a/b}:{\rm ker}^{+}S_{a,b}\to{\rm ker}^{-}S_{a,b},\qquad\phi_{+}\mapsto-\frac{a}{b}\phi_{+}$
(this is well-defined under the standing assumption (1.3)). In general we omit the index $a/b$ and write simply $M=M_{-a/b}$. So for $\phi_{+}\in H^{2}_{+}$ we have
(2.4) $\phi_{-}=M\phi_{+}\iff\phi_{-}\in H^{2}_{-}\quad\hbox{and}\quad a\phi_{+}+b\phi_{-}=0.$
Although $\mathop{\rm ker}\nolimits S_{a,b}$ is naturally a closed subspace of $L^{2}$, its projections ${\rm ker}^{+}S_{a,b}$ and ${\rm ker}^{-}S_{a,b}$ need not be. Further details are given in [11].
We now proceed to a discussion of near invariance.
Definition 2.1.
Let $\mathcal{M}$ be a subspace of $H^{2}_{+}$. Then we say that $\mathcal{M}$ is nearly invariant or nearly $S^{*}$-invariant if and only if
(2.5) $f\in\mathcal{M},\ f(0)=0\implies S^{*}f\in\mathcal{M}.$
We see that (2.5) is equivalent to the condition
$f\in\mathcal{M},\quad\overline{z}f\in H^{2}_{+}\implies\overline{z}f\in\mathcal{M}.$
It is well known that kernels of Toeplitz operators are nearly invariant, since if $\phi\in\mathop{\rm ker}\nolimits T_{g}$, so that $g\phi=\psi_{-}$ for some $\psi_{-}\in H^{2}_{-}$, and $\overline{z}\phi\in H^{2}_{+}$ we have $g\overline{z}\phi=\overline{z}\psi_{-}\in H^{2}_{-}$ and so $\overline{z}\phi\in\mathop{\rm ker}\nolimits T_{g}$.
Similarly ${\rm ker}^{+}S_{a,b}$, is nearly invariant, since if $\phi_{+}\in{\rm ker}^{+}S_{a,b}$ then there is a $\phi_{-}\in H^{2}_{-}$ such that $a\phi_{+}+b\phi_{-}=0$. Now if $\overline{z}\phi_{+}$ in $H^{2}_{+}$, then the identity $a(\overline{z}\phi_{+})+b\overline{z}\phi_{-}=0$ shows that $\overline{z}\phi_{+}\in{\rm ker}^{+}S_{a,b}$, since $\overline{z}\phi_{-}\in H^{2}_{-}$.
Now, although ${\rm ker}^{+}S_{a,b}$ need not be a closed subspace (see Example 3.10 below), it is known that the closure of a nearly invariant subspace is still nearly invariant [11, Prop. 4.4]. Closed nearly invariant spaces have been considered by many mathematicians, beginning with the work of Hitt [20] and Hayashi [19].
Let us now broaden the discussion by letting $\theta\in H^{\infty}$ be a non-constant inner function.
Definition 2.2.
A subspace $\mathcal{M}\subset H^{2}_{+}$ is nearly $\overline{\theta}$-invariant if $f\in\mathcal{M}$, $\overline{\theta}f\in H^{2}_{+}$ imply that $\overline{\theta}f\in\mathcal{M}$.
An argument similar to that above, replacing $\overline{z}$ by $\overline{\theta}$ shows that projected paired kernels ${\rm ker}^{+}S_{a,b}$ are nearly $\overline{\theta}$-invariant.
Note that $\phi_{+}\in{\rm ker}^{+}S_{a,b}\cap\theta H^{2}_{+}$ if and only if we can write $\phi_{+}=\theta\psi_{+}$ with $\psi_{+}\in H^{2}_{+}$ and $a\theta\psi_{+}+b\psi_{-}=0$ for some $\psi_{-}\in H^{2}_{-}$. Thus we conclude that
(2.6) ${\rm ker}^{+}S_{a,b}\cap\theta H^{2}_{+}=\theta\,{\rm ker}^{+}S_{a\theta,b}.$
Equivalently, we may write this as
(2.7) $\phi_{+}\in{\rm ker}^{+}S_{a,b},\quad\overline{\theta}\phi_{+}\in H^{2}_{+}\iff\phi_{+}\in\theta\,{\rm ker}^{+}S_{a\theta,b},$
and, thus, saying that ${\rm ker}^{+}S_{a,b}$ is nearly $\overline{\theta}$-invariant is equivalent to saying that
$\theta\,{\rm ker}^{+}S_{a\theta,b}\subset{\rm ker}^{+}S_{a,b}.$
By similar arguments we have also
(2.8) ${\rm ker}^{-}S_{a,b}\cap\overline{\theta}H^{2}_{-}=\overline{\theta}\,{\rm ker}^{-}S_{a,\overline{\theta}b},$
and so
(2.9) $\phi_{-}\in{\rm ker}^{-}S_{a,b},\quad\theta\phi_{-}\in H^{2}_{-}\iff\phi_{-}\in\overline{\theta}\,{\rm ker}^{-}S_{a,\overline{\theta}b}.$
In particular,
(2.10) ${\rm ker}^{+}S_{a,b}\cap zH^{2}_{+}=z{\rm ker}^{+}S_{az,b},$
consisting of the elements of ${\rm ker}^{+}S_{a,b}$ that vanish at 00; analogously, $\overline{z}{\rm ker}^{-}S_{a,\overline{z}b}$ consists of the functions $\phi_{-}\in{\rm ker}^{-}S_{a,b}$ such that $z\phi_{-}$ vanishes at $\infty$.
Also note that, for $\phi_{-}=M\phi_{+}$ (i.e., such that $a\phi_{+}+b\phi_{-}=0$ as in (2.3)), we have
(2.11) $\phi_{+}\in\theta\,{\rm ker}^{+}S_{a\theta,b}\iff\phi_{-}\in{\rm ker}^{-}S_{a,\overline{\theta}b}$
and
(2.12) $\phi_{+}\in{\rm ker}^{+}S_{a\theta,b}\iff\phi_{-}\in\overline{\theta}\,{\rm ker}^{-}S_{a,\overline{\theta}b}.$
3. Inclusion relations
Let us now study the relations between two paired kernels with symbol pairs that are related by multiplication operators, by means of certain inclusion relations between their projections into $H^{2}_{\pm}$. Although some of these relations were obtained in [11] we present here all proofs for completeness, which also allows us in some cases to present alternative proofs. We again remind the reader of our standing assumptions (1.3) and (1.4).
Proposition 3.1.
Let $\theta_{1}$ and $\theta_{2}$ be inner functions. Then
(3.1) ${\rm ker}^{+}S_{a\theta_{1},b\overline{\theta}_{2}}\subseteq{\rm ker}^{+}S_{a,b},$
and, for ${\rm ker}^{+}S_{a,b}\neq\{0\}$, the inclusion is proper if and only if $\theta_{1}\theta_{2}\not\in\mathbb{C}$; that is, $\theta_{1}$ or $\theta_{2}$ is non-constant.
Proof.
Let $\theta=\theta_{1}\theta_{2}$; then ${\rm ker}^{+}S_{a\theta_{1},b\overline{\theta_{2}}}={\rm ker}^{+}S_{a\theta,b}$. Let $\phi_{+}\in{\rm ker}^{+}S_{a\theta,b}$, so that $a\theta\phi_{+}+b\psi_{-}=0$ for some $\psi_{-}\in H^{2}_{-}$. Then we have $a\phi_{+}+b(\overline{\theta}\psi_{-})=0$ and so $\phi_{+}\in{\rm ker}^{+}S_{a,b}$.
If $\theta\in\mathbb{C}$, then the inclusion (3.1) is an equality. Otherwise, if $\theta\not\in\mathbb{C}$, then there exists $\phi_{-}\in{\rm ker}^{-}S_{a,b}$ such that $\phi_{-}\not\in\overline{\theta}H^{2}_{-}$, by near invariance: see Propositions 4.1 and 4.5 in [11].
Since $\phi_{-}\in{\rm ker}^{-}S_{a,b}$ but $\phi_{-}\not\in\overline{\theta}{\rm ker}^{-}S_{a,\overline{\theta}b}$ by (2.8), for $\phi_{+}=M^{-1}\phi_{-}$ we must have that $\phi_{+}\in{\rm ker}^{+}S_{a,b}$ but $\phi_{+}\not\in{\rm ker}^{+}S_{a\theta,b}$, by (2.12), so the inclusion is strict. ∎
By (2.1) we have the following.
Corollary 3.2.
Let $\theta_{1}$ and $\theta_{2}$ be inner functions. Then
(3.2) ${\rm ker}^{-}S_{a\theta_{1},b\overline{\theta}_{2}}\subseteq{\rm ker}^{-}S_{a,b},$
and, for ${\rm ker}^{-}S_{a,b}\neq\{0\}$, the inclusion is strict if and only if $\theta_{1}\theta_{2}\not\in\mathbb{C}$.
Proposition 3.3.
Let $h_{-}\in\overline{H^{\infty}}$. Then we have:
(i) ${\rm ker}^{+}S_{a,bh_{-}}\subseteq{\rm ker}^{+}S_{a,b}\subseteq{\rm ker}^{+}S_{ah_{-},b}$;
(ii) If $\overline{h_{-}}$ is outer then
(3.3) ${\rm ker}^{+}S_{a,bh_{-}}={\rm ker}^{+}S_{a,b}\quad\hbox{if}\quad\frac{a}{bh_{-}}\in L^{\infty},$
and
(3.4) ${\rm ker}^{+}S_{ah_{-},b}={\rm ker}^{+}S_{a,b}=\mathop{\rm ker}\nolimits T_{a/b}\quad\hbox{if}\quad a/b\in L^{\infty};$
(iii) If $\overline{h}_{-}$ is not outer, then the inclusions in (i) are strict.
Proof.
The inclusions in (i) are immediate consequences of the definition of a projected paired kernel. Regarding (ii), we have that if $a\phi_{+}+b\phi_{-}=0$ with $\phi_{\pm}\in H^{2}_{\pm}$, then
$a\phi_{+}+bh_{-}(h_{-}^{-1}\phi_{-})=0\quad\hbox{with}\quad h_{-}^{-1}\phi_{-}=-\frac{a}{bh_{-}}\phi_{+}\in L^{2}\cap\overline{z}\overline{\mathcal{N}_{+}},$
where $\mathcal{N}_{+}$ denotes the Smirnov class (see, e.g., [23]), if $\dfrac{a}{bh_{-}}\in L^{\infty}$. So $h_{-}^{-1}\phi_{-}\in H^{2}_{-}$, and therefore $\phi_{+}\in{\rm ker}^{+}S_{a,bh_{-}}$.
This, together with the first inclusion in (i), proves the first equality in (ii), and the second equality is proved analogously.
As for (iii), if $\overline{h_{-}}$ has a non-constant inner factor $\theta$, then it follows from Prop. 3.1 and from (i) that, if $h_{-}=\overline{\theta}\tilde{h}_{-}$ with $\tilde{h}_{-}\in\overline{H^{\infty}}$, then
${\rm ker}^{-}S_{a,b\overline{\theta}\tilde{h}_{-}}\subsetneq{\rm ker}^{+}S_{a,b\tilde{h}_{-}}\subseteq{\rm ker}^{+}S_{a,b},$
and analogously for the second inclusion. ∎
Remark 3.4.
Note that the inclusions in Prop. 3.3 (i) may be strict even if $\overline{h_{-}}$ is outer: see Example 3.10 below. Also note that from (3.4) we have that, if $g\in L^{\infty}$, $\overline{h_{-}}$ is outer and $h_{+}\in\mathcal{G}H^{\infty}$, then
(3.5) ${\rm ker}^{+}S_{gh_{-},h_{+}}=\mathop{\rm ker}\nolimits T_{gh_{+}^{-1}}=h_{+}\mathop{\rm ker}\nolimits T_{g}.$
Corollary 3.5.
Let $z_{0}\in\mathbb{T}$. Then
(i) if $a/b\in L^{\infty}$, then ${\rm ker}^{+}S_{a(z-z_{0}),b}={\rm ker}^{+}S_{az,b}$;
(ii) if $\dfrac{a}{b(z-z_{0})}\in L^{\infty}$, then ${\rm ker}^{+}S_{a,b(z-z_{0})}={\rm ker}^{+}S_{a,bz}$.
Proof.
Take $h_{-}=(z-z_{0})/z\in\overline{H^{\infty}}$, which is such that $\overline{h_{-}}$ is outer in $H^{\infty}$. Then (i) and (ii) follow from the second equality and first equality in Prop. 3.3 (ii), respectively. ∎
Note that
(3.6) ${\rm ker}^{+}S_{a,b}=\mathop{\rm ker}\nolimits T_{a/b},\quad\hbox{if}\quad a/b\in L^{\infty},$
so we have
$\mathop{\rm ker}\nolimits T_{g(z-z_{0})}=\mathop{\rm ker}\nolimits T_{gz}\quad\hbox{for}\quad z_{0}\in\mathbb{T}\quad\hbox{and}\quad g\in L^{\infty}.$
Taking (2.1) into account, we also have:
Corollary 3.6.
Let $h_{+}\in H^{\infty}$; then
${\rm ker}^{-}S_{ah_{+},b}\subseteq{\rm ker}^{-}S_{a,b}\subseteq{\rm ker}^{-}S_{a,bh_{+}}.$
Proposition 3.7.
Let $h_{+}\in H^{\infty}$. Then we have:
(i) $h_{+}{\rm ker}^{+}S_{ah_{+},b}\subseteq{\rm ker}^{+}S_{a,b}$ and $h_{+}{\rm ker}^{+}S_{a,b}\subseteq{\rm ker}^{+}S_{a,bh_{+}}$;
(ii) if $h_{+}\in\mathcal{G}H^{\infty}$, or if $h_{+}$ is outer with $\dfrac{b}{ah_{+}}\in L^{\infty}$, then
$h_{+}{\rm ker}^{+}S_{ah_{+},b}={\rm ker}^{+}S_{a,b};$
(iii) if $h_{+}\in\mathcal{G}H^{\infty}$, or if $h_{+}$ is outer with $\dfrac{b}{a}\in L^{\infty}$, then
$h_{+}{\rm ker}^{+}S_{a,b}={\rm ker}^{+}S_{a,bh_{+}};$
(iv) if $h_{+}$ is not outer, then the inclusions in (i) are strict.
Proof.
(i) The first inclusion is obvious, since for $\phi_{-}\in H^{2}_{-}$
$(ah_{+})\phi_{+}+b\phi_{-}=0\iff a(h_{+}\phi_{+})+b\phi_{-}=0.$
Regarding the second inclusion, if $\phi_{+}\in{\rm ker}^{+}S_{a,b}$ then $a(h_{+}\phi_{+})+(bh_{+})\phi_{-}=0$, so $h_{+}\phi_{+}\in{\rm ker}^{+}S_{a,bh_{+}}$.
(ii) It is left to show that ${\rm ker}^{+}S_{a,b}\subseteq h_{+}{\rm ker}^{+}S_{ah_{+},b}$. We have that
$a\phi_{+}+b\phi_{-}=0\iff(ah_{+})\frac{\phi_{+}}{h_{+}}+b\phi_{-}=0\iff\frac{\phi_{+}}{h_{+}}=-\frac{b}{ah_{+}}\phi_{-}.$
If $h_{+}\in\mathcal{G}H^{\infty}$, then ${\phi_{+}}/{h_{+}}\in H^{2}_{+}$; if $h_{+}$ is outer and $\dfrac{b}{ah_{+}}\in L^{\infty}$, then $\phi_{+}/h_{+}\in L^{2}\cap\mathcal{N}_{+}$ and so $\phi_{+}/h_{+}\in H^{2}_{+}$. It follows that $\phi_{+}/h_{+}\in{\rm ker}^{+}S_{ah_{+},b}$.
(iii) Now it remains to show that ${\rm ker}^{+}S_{a,bh_{+}}\subseteq h_{+}{\rm ker}^{+}S_{a,b}$. We have
$a\phi_{+}+(bh_{+})\phi_{-}=0\iff a\phi_{+}/h_{+}+b\phi_{-}=0.$
If $h_{+}\in\mathcal{G}H^{\infty}$, then $\phi_{+}/h_{+}\in H^{2}_{+}$; if $h_{+}$ is outer and $b/a\in L^{\infty}$ then $\phi_{+}/h_{+}\in L^{2}\cap\mathcal{N}_{+}=H^{2}_{+}$, so $\phi_{+}\in h_{+}{\rm ker}^{+}S_{a,b}$.
(iv) If $h_{+}$ is not outer, we cannot have ${\rm ker}^{+}S_{a,b}\subseteq h_{+}{\rm ker}^{+}S_{ah_{+},b}\subseteq h_{+}H^{2}_{+}$, nor ${\rm ker}^{+}S_{a,bh_{+}}\subseteq h_{+}{\rm ker}^{+}S_{a,b}\subseteq h_{+}H^{2}_{+}$, by near invariance (Cor. 4.2 in [11]). ∎
Remark 3.8.
Note that the inclusion in (i) above may be strict even if $h_{+}$ is outer (see also Remark 3.4).
Recalling (3.6), it follows from the second relation in Prop. 3.7 (i) that
(3.7) $h_{+}\mathop{\rm ker}\nolimits T_{a}\subseteq{\rm ker}^{+}S_{a,h_{+}}$
for any $h_{+}\in H^{\infty}$. In particular, since $\mathop{\rm ker}\nolimits T_{a}\neq\{0\}$ when $a\in\overline{H^{\infty}}$ and $\overline{a}$ is not outer, it follows from (3.7) that if $h_{+}$ or $\overline{h_{-}}$ is not outer, then
(3.8) ${\rm ker}^{+}S_{h_{-},h_{+}}\neq\{0\}.$
Corollary 3.9.
Let $h_{-}\in\overline{H^{\infty}}$. Then
(3.9) $h_{-}{\rm ker}^{-}S_{a,bh_{-}}\subseteq{\rm ker}^{-}S_{a,b}\quad\hbox{and}\quad h_{-}{\rm ker}^{-}S_{a,b}\subseteq{\rm ker}^{-}S_{ah_{-},b}.$
These inclusions are equalities if $\overline{h_{-}}\in\mathcal{G}H^{\infty}$ or $\overline{h_{-}}$ is outer, $\dfrac{a}{bh_{-}}\in L^{\infty}$; they are strict if $\overline{h_{-}}$ is not outer.
In the following example we illustrate some of the inclusion relations presented above.
Example 3.10.
Let $\theta$ be the singular inner function defined by
(3.10) $\theta(z)=\exp\left(\frac{z-1}{z+1}\right),\qquad(z\in\mathbb{T}).$
Then, by Prop. 3.7 (ii),
(3.11) $(z+1)K_{\theta}=(z+1){\rm ker}^{+}S_{\overline{\theta},1}={\rm ker}^{+}S_{\overline{\theta},z+1}.$
By Prop. 3.3 (i) and (3.5),
(3.12) $(z+1)K_{\theta}={\rm ker}^{+}S_{\overline{\theta}\overline{z},(z+1)/z}\subseteq{\rm ker}^{+}S_{\overline{\theta}\overline{z},1}=\mathop{\rm ker}\nolimits T_{\overline{\theta}\overline{z}}=K_{\theta z},$
where the inclusion is strict because, as shown in [21], $(z+1)K_{\theta}$ is not a closed subspace so, in particular, it cannot be a Toeplitz kernel. We thus see that the inclusion in Prop. 3.3 (i) may be strict if $\overline{h_{-}}$ is outer in $H^{\infty}$ but the condition $\dfrac{a}{bh_{-}}\in L^{\infty}$ is not satisfied, as happens in this case.
Remark 3.11.
Note that any space $uK_{\theta}\subset H^{2}_{+}$ with $u\in H^{2}_{+}$ outer, as in the example above, can be expressed as ${\rm ker}^{+}S_{a,b}$ if we allow for symbol pairs in $(L^{2})^{2}$, with the domain $\mathcal{D}=\{\phi\in L^{2}:a\phi_{+}+b\phi_{-}\in L^{2}\}$. Indeed, following the same reasoning as in Prop. 3.7 (ii), we have that $uK_{\theta}={\rm ker}^{+}S_{\overline{\theta},u}$.
On the other hand, if $u$ is not outer, then $uK_{\theta}$ cannot be expressed as a projected paired kernel, even if we allow for $L^{2}$ symbols, since this space cannot be contained in $\alpha H^{2}_{+}$ if $\alpha$ is a non-constant inner function, by an argument similar to that of Corollary 4.2 in [11].
We shall be interested in studying projected paired kernels of the form
(3.13) ${\rm ker}^{+}S_{\alpha B_{-},B_{+}}\quad\hbox{with}\quad B_{+},\overline{B_{-}}\in H^{\infty}.$
These are strictly contained in ${\rm ker}^{+}S_{B_{-},B_{+}}$, where we assume that either $\overline{B_{-}}$ or $B_{+}$ is not outer, so that ${\rm ker}^{+}S_{B_{-},B_{+}}\neq\{0\}$ by (3.8). In particular, we shall be interested in establishing conditions for (3.13) to be $\{0\}$.
It is thus convenient to see what form the previous inclusions take when $h_{+}$ or $\overline{h_{-}}$ are inner functions.
Proposition 3.12.
Let $\theta_{1}$ and $\theta$ be non-constant inner functions such that $\theta_{1}\prec\theta$ (that is, $\theta_{1}$ properly divides $\theta$). Then,
(i) ${\rm ker}^{+}S_{a\theta,b}\subsetneq{\rm ker}^{+}S_{a\theta_{1},b}\subsetneq{\rm ker}^{+}S_{a,b}\subsetneq{\rm ker}^{+}S_{a\overline{\theta_{1}},b}\subsetneq{\rm ker}^{+}S_{a\overline{\theta},b}$;
(ii) $\theta{\rm ker}^{+}S_{a\theta,b}\subsetneq\theta_{1}{\rm ker}^{+}S_{a\theta_{1},b}\subsetneq{\rm ker}^{+}S_{a,b}$,
assuming that the right-hand side of each inclusion is different from $\{0\}$.
Proof.
(i) follows from Prop. 3.1 and (ii) follows from Prop. 3.7. ∎
Analogously, we have:
Proposition 3.13.
Let $\theta_{1}$ and $\theta$ be non-constant inner functions such that $\theta_{1}\prec\theta$. Then
(i) ${\rm ker}^{-}S_{a,\overline{\theta}b}\subsetneq{\rm ker}^{-}S_{a,\overline{\theta_{1}}b}\subsetneq{\rm ker}^{-}S_{a,b}\subsetneq{\rm ker}^{-}S_{a,b\theta_{1}}\subset{\rm ker}^{-}S_{a,\theta b}$;
(ii) $\overline{\theta}{\rm ker}^{-}S_{a,\overline{\theta}b}\subsetneq\overline{\theta_{1}}{\rm ker}^{-}S_{a,{\overline{\theta}_{1}}b}\subsetneq{\rm ker}^{-}S_{a,b}$,
assuming that the right-hand side of each inclusion is different from $\{0\}$.
These inclusion relations allow us to answer practically, in some sense, the question of how much “smaller” ${\rm ker}^{+}S_{a\theta,b}$ or $\theta{\rm ker}^{+}S_{a\theta,b}$ are than ${\rm ker}^{+}S_{a,b}$ – a question which naturally arises from the strict inclusions of Proposition 3.1, and, in an analogous form, from Corollary 3.2, assuming that $\theta$ is a non-constant. Indeed, if $\theta_{0}=\theta$ and $\theta_{i}$ $(i=1,2,\ldots)$ are non-constant inner functions such that $\theta_{0}\succ\theta_{1}\succ\theta_{2}\succ\ldots$ and ${\rm ker}^{+}S_{a\theta,b}\neq\{0\}$, then we have
(3.14) ${\rm ker}^{+}S_{a\theta,b}\subsetneq{\rm ker}^{+}S_{a\theta_{1},b}\subsetneq{\rm ker}^{+}S_{a\theta_{2},b}\subsetneq\ldots\subsetneq{\rm ker}^{+}S_{a,b},$
and
(3.15) $\theta{\rm ker}^{+}S_{a\theta,b}\subsetneq\theta_{1}{\rm ker}^{+}S_{a\theta_{1},b}\subsetneq\theta_{2}{\rm ker}^{+}S_{a\theta_{2},b}\subsetneq\ldots\subsetneq{\rm ker}^{+}S_{a,b},$
which, according to (2.11) and (2.12), are equivalent to
(3.16) $\overline{\theta}{\rm ker}^{-}S_{a,\overline{\theta}b}\subsetneq\theta_{1}{\rm ker}^{-}S_{a,\overline{\theta_{1}}b}\subsetneq\theta_{2}{\rm ker}^{-}S_{a,\overline{\theta_{2}}b}\subsetneq\ldots\subseteq{\rm ker}^{-}S_{a,b},$
and
(3.17) ${\rm ker}^{-}S_{a,\overline{\theta}b}\subsetneq{\rm ker}^{-}S_{a,\overline{\theta_{1}}b}\subsetneq{\rm ker}^{-}S_{a,\overline{\theta_{2}}b}\subsetneq\ldots\subsetneq{\rm ker}^{-}S_{a,b}.$
If $\theta$ is a finite Blaschke product, then we have the following.
Proposition 3.14.
[11, Thm. 6.8] Let $B$ be a finite Blaschke product. Then
(i) ${\rm ker}^{+}S_{aB,b}$ is finite-dimensional if and only if ${\rm ker}^{+}S_{a,b}$ is finite-dimensional, and analogously for ${\rm ker}^{+}S_{a,Bb}$;
(ii) if $\dim{\rm ker}^{+}S_{a,b}=d<\infty$ and $\dim K_{B}\geq d$, then ${\rm ker}^{+}S_{aB,b}=\{0\}$.
(iii) if $\dim{\rm ker}^{+}S_{a,b}=d<\infty$ and $\dim K_{B}=k<d$, then $\dim{\rm ker}^{+}S_{aB,b}=d-k$.
From (3.14) and (3.15) we also have:
Proposition 3.15.
Let $B=B_{0}$ be a finite Blaschke product of degree $k$, and let $B_{1},B_{2},\ldots,B_{k}$ be finite Blaschke products such that $B\succ B_{1}\succ B_{2}\succ\ldots\succ B_{k}\in\mathbb{C}$. Moreover, suppose that $\dim{\rm ker}^{+}S_{a,b}>k$. Then there exist $\phi_{i+}\in{\rm ker}^{+}S_{aB_{i},b}\setminus{\rm ker}^{+}S_{aB_{i-1}b}$, for $i=1,2,\ldots,k$, such that
(3.18) ${\rm ker}^{+}S_{a,b}={\rm ker}^{+}S_{aB,b}+\mathop{\rm span}\nolimits\{\phi_{1+},\ldots,\phi_{k+}\}.$
and there exist $\psi_{1+}\in B_{i}{\rm ker}^{+}S_{aB_{i}}\setminus B_{i-1}{\rm ker}^{+}S_{a,B_{i-1},b}$, for $i=1,2,\ldots,k$, such that
(3.19) ${\rm ker}^{+}S_{a,b}=B{\rm ker}^{+}S_{aB,b}+\mathop{\rm span}\nolimits\{\psi_{1+},\ldots,\psi_{k+}\}.$
Proof.
It is enough to prove the result when $B$ is a finite Blaschke product of degree $1$. Let us take $B=z$ for simplicity: the general argument is similar. Since $z{\rm ker}^{+}S_{az,b}\subsetneq{\rm ker}^{+}S_{a,b}$, there exists $\widetilde{\psi}_{+}\in{\rm ker}^{+}S_{a,b}\setminus z{\rm ker}^{+}S_{az,b}$, so $\widetilde{\psi}_{+}(0)\neq 0$. Now for $\phi_{+}\in{\rm ker}^{+}S_{a,b}$ define
$\widetilde{\phi}_{+}=\widetilde{\psi}_{+}(0)\phi_{+}-\phi_{+}(0)\widetilde{\psi}_{+}.$
We have that
$\widetilde{\phi}_{+}\in{\rm ker}^{+}S_{a,b}\cap zH^{2}_{+}=z{\rm ker}^{+}S_{az,b}$
and
$\phi_{+}=\underbrace{\frac{\widetilde{\phi}_{+}}{\widetilde{\psi}_{+}(0)}}_{\in\,z{\rm ker}^{+}S_{az,b}}+\underbrace{\frac{\phi_{+}(0)}{\widetilde{\psi}_{+}(0)}\widetilde{\psi}_{+}}_{\in\,{\rm ker}^{+}S_{a,b}\setminus z{\rm ker}^{+}S_{az,b}},$
so we conclude that (3.18) holds for $B=z$.
By (2.11) it follows from this that
${\rm ker}^{-}S_{a,b}={\rm ker}^{-}S_{a,\overline{z}b}+\mathop{\rm span}\nolimits\{\widetilde{\phi}_{-}\},$
where $\widetilde{\phi}_{-}=M\widetilde{\phi}_{+}\in{\rm ker}^{-}S_{a,b}\setminus{\rm ker}^{-}S_{a,\overline{z}b}$. Now, using (2.2) we conclude that
${\rm ker}^{+}S_{a,b}={\rm ker}^{+}S_{az,b}+\mathop{\rm span}\nolimits\{\phi_{+}\},$
where $\phi_{+}\in{\rm ker}^{+}S_{a,b}\setminus{\rm ker}^{+}S_{az,b}$. ∎
Finally, note that for a projected paired kernel of the form (3.13), if $B_{-}=\overline{\theta_{1}}\overline{O_{1}}$ with $\theta_{1}$ inner and $O_{1}$ outer in $H^{\infty}$, and $B_{+}=\theta_{2}O_{2}$ with $\theta_{2}$ inner and $O_{2}\in\mathcal{G}H^{\infty}$, then by (3.5) we have that
(3.20) ${\rm ker}^{+}S_{\alpha B_{-},B_{+}}={\rm ker}^{+}S_{\alpha\overline{\theta_{1}}\overline{\theta_{2}}\overline{O_{1}},O_{2}}=O_{2}\mathop{\rm ker}\nolimits T_{\alpha\overline{\theta}}$
with $\theta=\theta_{1}\theta_{2}$, so the question whether ${\rm ker}^{+}S_{\alpha B_{-},B_{+}}=\{0\}$ reduces to the question whether $\mathop{\rm ker}\nolimits T_{\alpha\overline{\theta}}=\{0\}$. This, in turn, is related to the question of existence of a multiplier from $K_{\theta}$ into $K_{\alpha}$, studied in [14, 9]. The answer to our question, in the case when one of the inner functions is a finite Blaschke product, was given in [6, Sec. 6], but the situation is more complicated when neither inner function is a finite Blaschke product.
So suppose that $\theta$ and $\alpha$ are non-constant inner functions in $H^{\infty}(\mathbb{D})$ with no nontrivial common factors. The question is: when is $\mathop{\rm ker}\nolimits T_{\alpha\overline{\theta}}$ different from $\{0\}$? By Coburn’s lemma at least one of $\mathop{\rm ker}\nolimits T_{\alpha\overline{\theta}}$ and $\mathop{\rm ker}\nolimits T_{\theta\overline{\alpha}}$ will be zero.
Recall that the spectrum of an inner function $\sigma(\theta)$ consists of all $\lambda\in\overline{\mathbb{D}}$ such that $\liminf_{z\to\lambda}|\theta(z)|=0$, and it is the union of the closure of its zero set in $\mathbb{D}$ and the support of its singular measure (if any). Useful facts from [15, Sec. 7.3] are:
  • if $\sigma(\theta)\cap\mathbb{T}\neq\mathbb{T}$, then $\theta$ has an analytic continuation across $\mathbb{T}\setminus\sigma(\theta)$;
  • $\theta$ does not have an analytic continuation across any point of $\sigma(\theta)\cap\mathbb{T}$;
  • every function in $K_{\theta}$ has an analytic continuation across $\mathbb{T}\setminus\sigma(\theta)$.
Theorem 3.16.
If $\mathop{\rm ker}\nolimits T_{\alpha\overline{\theta}}\neq\{0\}$ we have $\sigma(\alpha)\cap\mathbb{T}\subset\sigma(\theta)\cap\mathbb{T}$.
Proof.
We may suppose that $\sigma(\theta)\cap\mathbb{T}\neq\mathbb{T}$, that $f_{+}\neq 0$ and $f_{+}\in\mathop{\rm ker}\nolimits T_{\overline{\theta}}$. That is, $\alpha\overline{\theta}f_{+}=f_{-}$ for some $f_{-}\in H^{2}_{-}$. Then $f_{+}$ and $\alpha f_{+}$ both lie in $K_{\theta}$, and have analytic continuations across $\mathbb{T}\setminus\sigma(\theta)$. Hence, by division, $\alpha$ has a meromorphic continuation, but since it is inner, this must be an analytic continuation (no poles), since it has to be unimodular a.e. on $\mathbb{T}$. ∎
Note that if $\theta=\alpha\beta$ with $\beta$ inner, non-constant, then $\mathop{\rm ker}\nolimits T_{\overline{\theta}\alpha}=K_{\beta}\neq\{0\}$, and we do indeed have $\sigma(\alpha)\subset\sigma(\theta)$. Moreover, the example given in [6, Ex. 6.3] corresponds to $\theta$ and $\alpha$ being singular inner functions with point masses at different points, so we recover the result that $\mathop{\rm ker}\nolimits T_{\overline{\theta}\alpha}=\{0\}$ in this case. This leaves open many other examples, e.g., what happens if $\theta$ and $\alpha$ are Blaschke products with disjoint zero sets that both converge to $1$?
4. Kernels of ATTO
We defined asymmetric truncated Toeplitz operators (ATTO) in (1.11) and now consider the question of describing the kernel of such an operator when the symbol has the form
(4.1) $\phi=\overline{\theta}\frac{A_{1-}}{B_{1-}}-\alpha\frac{A_{2+}}{B_{2+}}\in L^{\infty},$
with $(A_{2+},B_{2+})\in{\rm CP}^{+}$ and $(A_{1-},B_{1-})\in{\rm CP}^{-}$. Here ${\rm CP}^{+}$ denotes the set of all corona pairs $(f_{1+},f_{2+})$ with $f_{1+},f_{2+}\in H^{\infty}$ such that there exist $\tilde{f}_{1+},\tilde{f}_{2+}\in H^{\infty}$ satisfying
(4.2) $f_{1+}\tilde{f}_{1+}+f_{2+}\tilde{f}_{2+}=1.$
By the well-known corona theorem [12], there exist $\tilde{f}_{1+}$ and $\tilde{f}_{2+}$ such that (4.2) holds if and only if
(4.3) $\inf_{z\in\mathbb{D}}|f_{1+}(z)|+|f_{2+}(z)|>0.$
${\rm CP}^{-}$ is defined analogously, with $H^{\infty}$ replaced by $\overline{H^{\infty}}$.
This class of ATTO is motivated by the study of finite-rank ATTO, which will be considered as a particular case where $A_{1-},B_{1-}$ and $A_{2+},B_{2+}$ are polynomials (see [8, 11]. The starting point of this study is a result from [10], which is formulated here as follows, denoting by $\mathcal{F}$ the set of all complex-valued measurable functions defined almost everywhere on $\mathbb{T}$.
Theorem 4.1.
[10, Cor. 3.4] Let $f=(f_{1},f_{2})$ and $g=(g_{1},g_{2})$ belong to $\mathcal{F}^{2}$, with left inverses $\tilde{f}^{T}$ and $\tilde{g}^{T}$, respectively, where $\tilde{f}=(\tilde{f}_{1},\tilde{f}_{2})$ and $\tilde{g}=(\tilde{g}_{1},\tilde{g}_{2})$ belong to $\mathcal{F}^{2}$; i.e.,
(4.4) $\tilde{f}^{T}f=\tilde{g}^{T}g=1.$
Suppose also that, for a given $G\in(L^{\infty})^{2\times 2}$ we have that
(4.5) $Gf=g.$
Let moreover
(4.6) $\mathcal{S}=(\det G.f^{T}(H^{2}_{+})^{2})\cap(g^{T}(H^{2}_{-})^{2}),$
and
(4.7) $\mathcal{K}=(\tilde{f}^{T}\mathop{\rm ker}\nolimits H_{f\tilde{f}^{T}})\cap(\tilde{g}^{T}\mathop{\rm ker}\nolimits\widetilde{H}_{g\tilde{g}^{T}}),$
where
(4.8) $\mathop{\rm ker}\nolimits H_{f\tilde{f}^{T}}=\{\phi_{+}\in(H^{2}_{+})^{2}:P^{-}f\tilde{f}^{T}\phi_{+}=0\}$
and
(4.9) $\mathop{\rm ker}\nolimits\widetilde{H}_{g\tilde{g}^{T}}=\{\phi_{-}\in(H^{2}_{-})^{2}:P^{+}g\tilde{g}^{T}\phi_{-}=0\}.$
Then, if $\mathcal{S}=\{0\}$ we have that
(4.10) $\mathop{\rm ker}\nolimits T_{G}=\mathcal{K}f.$
Remark 4.2.
Given that $f$ is a fixed vector function and $\mathcal{K}$ is a space of scalar functions we say that $\mathop{\rm ker}\nolimits T_{G}$ is of scalar type if it satisfies an equation of the form (4.10).
There are many examples of scalar-type kernels for block Toeplitz operators. We have the following analogue of Coburn’s lemma for Toeplitz operators with $2\times 2$ symbols.
Theorem 4.3.
[10] If $G\in(L^{\infty})^{2\times 2}$ and $\det G\in L^{\infty}\setminus\{0\}$ then either $\mathop{\rm ker}\nolimits T_{G}=\{0\}$ or $\mathop{\rm ker}\nolimits T_{G}^{*}=\{0\}$, or both kernels are of scalar type.
Applying Theorem 4.1 to the study of kernels of ATTO will be based on the following result.
Theorem 4.4.
[8]. Let $\theta,\alpha$ be inner functions, let $\phi\in L^{\infty}$, and set
(4.11) $G=\begin{bmatrix}\overline{\theta}&0\\ \phi&\alpha\end{bmatrix}.$
Denoting by $P_{1}$ the projection $(x,y)\overset{P_{1}}{\longmapsto}x$, we have that
(4.12) $\mathop{\rm ker}\nolimits A^{\theta,\alpha}_{\phi}=P_{1}\mathop{\rm ker}\nolimits T_{G}.$
To study the kernel of an ATTO of the form $A_{\phi}^{\theta,\alpha}$, with $\phi$ given by (4.1), we shall therefore apply Theorem 4.1 to matrix symbols of the form (4.11). A particular type of triangular symbol in this class was also studied, in the context of almost-periodic factorization on the real line, in the book [2].
It is easy to see that (4.5) is satisfied for
(4.13) $f=\theta\begin{bmatrix}1\\ \frac{A_{2+}}{B_{2+}}\end{bmatrix},\qquad g=\begin{bmatrix}1\\ \frac{A_{1-}}{B_{1-}}\end{bmatrix},$
which have left inverses $\tilde{f}^{T}$ and $\tilde{g}^{T}$ respectively, with
(4.14) $\tilde{f}=\begin{bmatrix}\,\overline{\theta}\,\\ 0\end{bmatrix},\qquad\tilde{g}=\begin{bmatrix}1\\ 0\end{bmatrix}.$
It follows from Theorems 4.1 and 4.4 that, taking
(4.15) $G=\begin{bmatrix}\overline{\theta}&0\\ \overline{\theta}\frac{A_{1-}}{B_{1-}}-\alpha\frac{A_{2+}}{B_{2+}}&\alpha\end{bmatrix}$
and $f,g,\tilde{f},\tilde{g}$ as in (4.13) and (4.14), we have the following.
Theorem 4.5.
If
(4.16) $\mathcal{S}:=\left(\alpha\begin{bmatrix}1&\frac{A_{2+}}{B_{2+}}\end{bmatrix}(H^{2}_{+})^{2}\right)\cap\left(\begin{bmatrix}1&\frac{A_{1-}}{B_{1-}}\end{bmatrix}(H^{2}_{-})^{2}\right)=\{0\},$
then, for $\phi$ given by (4.1),
(4.17) $\mathop{\rm ker}\nolimits A^{\theta,\alpha}_{\phi}=\mathcal{K}\theta,$
where
(4.18) $\mathcal{K}=\biggl{(}\begin{bmatrix}\,\overline{\theta}&0\end{bmatrix}\underbrace{\mathop{\rm ker}\nolimits H_{\begin{bmatrix}1&0\\ \frac{A_{2+}}{B_{2+}}&0\end{bmatrix}}}_{\mathop{\rm ker}\nolimits H_{f\tilde{f}^{T}}}\biggr{)}\cap\biggl{(}\begin{bmatrix}1&0\end{bmatrix}\underbrace{\mathop{\rm ker}\nolimits\widetilde{H}_{\begin{bmatrix}1&0\\ \frac{A_{1-}}{B_{1-}}&0\end{bmatrix}}}_{\mathop{\rm ker}\nolimits\widetilde{H}_{g\tilde{g}^{T}}}\biggr{)}.$
Let us first show that, if $K_{\alpha}$ is “big” enough, then $\mathcal{S}=\{0\}$ independently from what the inner function $\theta$ is.
Proposition 4.6.
Let $\mathcal{S}$ be defined as in (4.16). Then
(4.19) $B_{2+}\mathcal{S}\subseteq\alpha{\rm ker}^{+}S_{\alpha B_{1-},B_{2+}}.$
Proof.
$\mathcal{S}$ consists of functions of the form
(4.20) $\alpha\left(\phi_{1+}+\frac{A_{2+}}{B_{2+}}\phi_{2+}\right)\qquad\hbox{with}\quad\phi_{1+},\phi_{2+}\in H^{2}_{+},$
(4.21) $\alpha\left(\phi_{1+}+\frac{A_{2+}}{B_{2+}}\phi_{2+}\right)=\phi_{1-}+\frac{A_{1-}}{B_{1-}}\phi_{2-}\qquad\hbox{for some}\quad\phi_{1-},\phi_{2-}\in H^{2}_{-},$
so
$B_{1-}B_{2+}\left(\alpha(\phi_{1+}+\frac{A_{2+}}{B_{2+}}\phi_{2+})\right)$ $=\alpha B_{1-}(B_{2+}\phi_{1+}+A_{2+}\phi_{2+})$
$=B_{2+}(B_{1-}\phi_{1-}+A_{1-}\phi_{2-}).$
Thus
$B_{1-}B_{2+}\mathcal{S}\subseteq(\alpha B_{1-}H^{2}_{+})\cap(B_{2+}H^{2}_{-}),$
which, using the lemma below, means that
$B_{1-}B_{2+}\mathcal{S}\subseteq\alpha B_{1-}{\rm ker}^{+}S_{\alpha B_{1-},B_{2+}}$
as required. ∎
At this point it will be helpful to observe that ${\rm ker}^{+}S_{A,-B}={\rm ker}^{+}S_{A,B}$ for all $A,B\in L^{\infty}$.
Lemma 4.7.
Let $A,B\in L^{\infty}$. Then $AH^{2}_{+}\cap BH^{2}_{-}=A\,{\rm ker}^{+}S_{A,B}$.
Proof.
If $A\phi_{+}=B\phi_{-}$, with $\phi_{\pm}\in H^{2}_{\pm}$, then $A\phi_{+}-B\phi_{-}=0$ and $\phi_{+}\in{\rm ker}^{+}S_{A,B}$, so $AH^{2}_{+}\cap BH^{2}_{-}\subseteq\,A\,{\rm ker}^{+}S_{A,B}$.
Conversely, it is clear that $A\,{\rm ker}^{+}S_{A,B}\subseteq AH^{2}_{+}$, and from $A\phi_{+}-B\phi_{-}=0$ we see that we also have $A\,{\rm ker}^{+}S_{A,B}\subseteq BH^{2}_{-}$. ∎
Corollary 4.8.
If ${\rm ker}^{+}S_{\alpha B_{1-},B_{2+}}=\{0\}$ then $\mathcal{S}=\{0\}$.
Now let $\overline{B_{1-}}=I_{1}O_{1}$ and $B_{2+}=I_{2}O_{2}$ be inner–outer factorizations, with $I_{1},I_{2}$ inner and $O_{1},O_{2}$ outer. Let $I=I_{1}I_{2}$.
Corollary 4.9.
If $O_{2}\in\mathcal{G}H^{\infty}$ then
(4.22) $\mathcal{S}=\{0\}\quad\hbox{if}\quad\mathop{\rm ker}\nolimits T_{\alpha\overline{I}}=\{0\},$
i.e., if $\alpha$ “annihilates” $K_{I}$ [6, Def. 6.1].
Proof.
This is an immediate consequence of Corollary 4.8 and (3.20). ∎
With the same notation, from the previous result and Proposition 3.14 we thus have:
Proposition 4.10.
If $I_{1}$ and $I_{2}$ are finite Blaschke products and $O_{2}\in\mathcal{G}H^{\infty}$, then $\mathcal{S}=\{0\}$ if
(4.23) $\dim K_{\alpha}\geq\dim K_{I_{1}}+\dim K_{I_{2}}.$
Now assume that $\mathcal{S}=\{0\}$. In that case, by Theorem 4.5 we have (4.17). To characterise $\mathcal{K}$, we will use the following result, with the notation
(4.24) $\mathcal{H}_{X}=\{h\in\mathcal{H}:Xh\in\mathcal{H}\}.$
Proposition 4.11.
Let $A_{2+},B_{2+}\in H^{\infty}$ with $\inf_{z\in\mathbb{D}}|A_{2+}(z)|+|B_{2+}(z)|>0$. Then
(4.25) $(H^{2}_{+})_{A_{2+}/B_{2+}}=\mathop{\rm ker}\nolimits H_{A_{2+}/B_{2+}}=B_{2+}H^{2}_{+}.$
Analogously, if $A_{1-},B_{1-}\in\overline{H^{\infty}}$ with $\inf_{z\in\mathbb{D}}|\overline{A_{1-}}(z)|+|\overline{B_{1-}}(z)|>0$, then
(4.26) $(H^{2}_{-})_{A_{1-}/B_{1-}}=\mathop{\rm ker}\nolimits\widetilde{H}_{A_{1-}/B_{1-}}=B_{1-}H^{2}_{-}.$
Proof.
Clearly, $B_{2+}H^{2}_{+}\subseteq(H^{2}_{+})_{A_{2+}/B_{2+}}$. Conversely, if $\dfrac{A_{2+}}{B_{2+}}\phi_{1+}=\psi_{+}$, with $\phi_{1+},\psi_{+}\in H^{2}_{+}$, and $A_{2+}\widetilde{A}_{2+}+B_{2+}\widetilde{B}_{2+}=1$, then
$\widetilde{A}_{2+}A_{2+}\phi_{1+}=\widetilde{A}_{2+}B_{2+}\psi_{+}$ $\iff\phi_{1+}-B_{2+}\widetilde{B}_{2_{+}}\phi_{1+}=\widetilde{A}_{2+}B_{2+}\psi_{+}$
$\iff\phi_{1+}=B_{2+}(\widetilde{A}_{2+}\psi_{+}+\widetilde{B}_{2+}\phi_{1+})\in B_{2+}H^{2}_{+}.$
The proof of (4.26) is analogous. ∎
Remark 4.12.
Note in (4.25) that $(H^{2}_{+})_{A_{2+}/B_{2+}}$ can be seen as the kernel of a possibly unbounded Hankel operator, and analogously for (4.26). Thus Proposition 4.11 establishes a link between the corona theorem and kernels of Hankel operators with symbols of a certain form.
Thus we have, using Proposition 4.11:
Proposition 4.13.
Let $f,g,\tilde{f},\tilde{g}$ be as defined in (4.13) and (4.14). Then
(4.27) $\mathop{\rm ker}\nolimits H_{f\tilde{f}^{T}}=(B_{2+}H^{2}_{+},H^{2}_{+})$
and
(4.28) $\mathop{\rm ker}\nolimits\widetilde{H}_{g\tilde{g}^{T}}=(B_{1-}H^{2}_{-},H^{2}_{-}).$
Finally, from Theorem 4.5 we get the following.
Theorem 4.14.
Let $\phi$ be defined as in (4.1) and suppose that $\mathcal{S}=\{0\}$. Then $\mathop{\rm ker}\nolimits A^{\theta,\alpha}_{\phi}$ does not depend on $\alpha$ and we have
(4.29) $\mathop{\rm ker}\nolimits A^{\theta,\alpha}_{\phi}=B_{2+}{\rm ker}^{+}S_{\overline{\theta}B_{2+},-B_{1-}}.$
Further, if $\overline{B}_{1-}=I_{1}O_{1}$ and $B_{2+}=I_{2}O_{2}$ are inner–outer factorizations, with $O_{1},O_{2}\in\mathcal{G}H^{\infty}$, then
(4.30) $\mathop{\rm ker}\nolimits A^{\theta,\alpha}_{\phi}=I_{2}\mathop{\rm ker}\nolimits T_{\overline{\theta}(I_{1}I_{2})}.$
Proof.
The first part follows from Theorem 4.5 and (4.30) follows from Propositions 3.3 and 3.7. ∎
Recall that the question whether Toeplitz kernels of the form arising in (4.30) are equal to $\{0\}$ was discussed in Section 3. We also note that the results here do not depend on $A_{1-}$ nor $A_{2+}$.
Finally, let us consider the particular case of ATTO of finite rank, studied in [11], where the symbol takes the form
(4.31) $\phi=\overline{\theta}R_{+}-\alpha R_{-}+\sum_{j=1}^{N}\frac{\overline{\theta}P^{\alpha}_{n_{j}-1}(z)-\alpha P^{\overline{\theta}}_{n_{j}-1}(t_{j})}{(z-t_{j})^{n_{j}}},$
where
  1. (i)
    $R_{\pm}$ are rational functions vanishing at $\infty$, such that $R_{+}\in H^{\infty}$ and $R_{-}\in\overline{H^{\infty}}$;
  2. (ii)
    $t_{j}\in\mathbb{T}$ for $j=1,2,\ldots,N$, and are regular points for $\theta$ and $\alpha$;
  3. (iii)
    $P^{\alpha}_{n_{j}-1}$ and $P^{\overline{\theta}}_{n_{j}-1}$ are the Taylor polynomials of order $n_{j}-1$ relative to the point $t_{j}$, for $\alpha$ and $\overline{\theta}$ respectively.
The formula in (4.31) can be rewritten in the form
(4.32) $\phi=\overline{\theta}\frac{Q_{1}}{\mathcal{E}D_{1+}}-\alpha\frac{Q_{2}}{\mathcal{E}D_{2-}}\in L^{\infty},$
where:
  • $Q_{1}$ and $Q_{2}$ are polynomials, sharing no common zeros with $\mathcal{E}D_{1+}$ and $\mathcal{E}D_{2-}$, respectively;
  • $D_{1+}$ is a polynomial of degree $N_{1}$ with all its zeros in $\mathbb{D}^{e}:=\{z\in\mathbb{C}:|z|>1\}$;
  • $D_{2-}$ is a polynomial of degree $N_{2}$ with all its zeros in $\mathbb{D}$;
  • $\mathcal{E}$ is a polynomial of degree $N_{0}$ with all its zeros on $\mathbb{T}$;
  • $\deg Q_{1}<N_{0}+N_{1}$ and $\deg Q_{2}<N_{0}+N_{2}$.
Therefore
(4.33) $\phi=\overline{\theta}\frac{A_{1-}}{B_{1-}}-\alpha\frac{A_{2+}}{B_{2+}},$
with
$\begin{array}[]{ll}A_{1-}=\frac{Q_{1}}{z^{N_{0}+N_{1}}}&B_{1-}=\frac{\mathcal{E}D_{1+}}{z^{N_{0}+N_{1}}}\\ \\ A_{2+}=Q_{2}&B_{2_{+}}=\mathcal{E}D_{2-}.\end{array}$
By Corollary 4.8 we know that $\mathcal{S}=\{0\}$ if ${\rm ker}^{+}S_{\alpha\mathcal{E}D_{1+}/z^{N_{0}+N_{1}},\mathcal{E}D_{2-}}=\{0\}$. Now,
${\rm ker}^{+}S_{\alpha\mathcal{E}D_{1+}/z^{N_{0}+N_{1}},\mathcal{E}D_{2-}}$ $={\rm ker}^{+}S_{\alpha D_{1+}/z^{N_{0}+N_{1}},D_{2-}}$
$={\rm ker}^{+}S_{\alpha\overline{z}^{N_{0}+N_{1}}D_{1+},z^{N_{2}}(D_{2-}/z^{N_{2}})}$
$=D_{1+}^{-1}{\rm ker}^{+}S_{\alpha\overline{z}^{N_{0}+N_{1}},z^{N_{2}}}$
$=D_{1+}^{-1}\mathop{\rm ker}\nolimits T_{\alpha\overline{z}^{N_{0}+N_{1}+N_{2}}}$
because $D_{1+}\in\mathcal{G}H^{\infty}$, $D_{2-}/z^{N_{2}}\in\mathcal{G}\overline{H^{\infty}}$ (by Propositions 3.3 and 3.7 and by (3.6)). Therefore
$\mathcal{S}=\{0\}\qquad\hbox{if}\quad\deg\alpha>N_{0}+N_{1}+N_{2}.$
Now, by Theorem 4.14,
$\mathop{\rm ker}\nolimits A_{\phi}^{\theta,\alpha}$ $=\mathcal{E}D_{2-}{\rm ker}^{+}S_{\overline{\theta}\mathcal{E}D_{2-},-\mathcal{E}D_{1+}/z^{N_{0}+N_{1}}}$
$=\mathcal{E}D_{2-}{\rm ker}^{+}S_{\overline{\theta}D_{2-},-D_{1+}/z^{N_{0}+N_{1}}}$
$=\mathcal{E}D_{2-}{\rm ker}^{+}S_{\overline{\theta}z^{N_{2}}(D_{2-}/z^{N_{2}})z^{N_{0}+N_{1}},-D_{1+}}$
$=\mathcal{E}D_{2-}D_{1+}{\rm ker}^{+}S_{\overline{\theta}z^{N_{0}+N_{1}+N_{2}},-1}$
$=\mathcal{E}D_{2-}D_{1+}\mathop{\rm ker}\nolimits T_{\overline{\theta}z^{N_{0}+N_{1}+N_{2}}}\,,$
by Propositions 3.3 and 3.7, as above. Thus we obtain the same result as in [11], but in a simpler way.
Acknowledgment
Work partially funded by FCT/Portugal through project UIDB/04459/2020 with DOI identifier 10-54499/UIDP/04459/2020